OpenStudy (anonymous):

solve the system of equations: 2x + 3y = 1 y= 3x + 15

OpenStudy (anonymous):

wow, that is great, I would have never figured that one out. Thanks so much. so x = 4 the answer or -7/3?

OpenStudy (anonymous):

the answer written out should be in ( , ) would it be like (-7, 3)

OpenStudy (anonymous):

x = 4 and y = -7/3, check if the answer is right for me? :P

OpenStudy (anonymous):

oh then its (4, -7/3) that is (x, y)

OpenStudy (anonymous):

this is very confusing. Thanks again.

OpenStudy (anonymous):

hang on.. its x = -4 and y = 3

OpenStudy (anonymous):

I have another that has been making me crazy it is (-3,2) with a slope of -2. what is the slope intercept?

OpenStudy (anonymous):

@rational what is slope intercept?

OpenStudy (anonymous):

2x + 3y = 1 y= 3x + 15 since y value is given, just plug that in the first equation 2x + 3(3x + 15) = 1 2x + 9x + 45 = 1 11x + 45 =1 11x = -44 x = -4 now plug x value, 2(-4) + 3y =1 -8 + 3y = 1 3y = 9 y = 3

OpenStudy (anonymous):

I'm so sorry, I got the wrong answer earlier, that's why I told you to check with your book

OpenStudy (anonymous):

The question is what is the slope intercept form of the function that contains the point (-3,2) and has a slope of -2? y= x+

OpenStudy (anonymous):

wait for @rational for this.. I'm not very sure

OpenStudy (anonymous):

or @dan815

OpenStudy (rational):

slope intercept form : \[y = mx + b\] where \(m\) = slope \(b\) = y intercept

OpenStudy (rational):

you're given slope = \(m=-2\) so \[y = -2x+b\] use the given point and find the value of \(b\)

OpenStudy (anonymous):

so where does -3 and 2 come in?

OpenStudy (anonymous):

OH I REMEMBER NOW! xD y = mx + b, that's my next chapter's pre-cursor.. tangent to normal, curve and stuff.. lovely, @rational is gonna be spammed with questions tomorrow :P

OpenStudy (anonymous):

ok, so would it be y=2 x + 3